Frege’s Contributions to Foundational Studies

Richard Heck

Gottlob Frege is well-known as the inventor of modern logic, in particular, of quantification theory, and as a defender of the view that arithmetic is “a branch of logic”. The details of his views, and of his formal work, are less well-known, however. I intend to provide an overview of these matters and then to discuss, with reference to some recent work by Jamie Tappenden and Patricia Blanchette, aspects of Frege’s famous dispute with Hilbert regarding proofs of the independence of geometrical axioms.

Address: 208 Emerson Hall, Harvard University, Cambridge MA 02138

E-mail: heck@fas.harvard.edu